### v1

# Sharp Estimates for the Robin Laplacian under a Perimeter Constraint in hyperbolic space

Abstract.
In this paper, we establish a lower bound, in terms of the isoperimetric deficit, for the first eigenvalue of the Robin Laplacian with negative boundary parameter on horospherically convex bounded domains in hyperbolic space, which implies that the geodesic ball maximizes this eigenvalue among all such domains. Furthermore, we derive upper bounds for the first eigenvalue of the Robin Laplacian with positive boundary parameter on horospherically convex bounded domains in hyperbolic space.
Key words and phrases:
Robin Laplacian; isoperimetric inequalities; horospherically convex; boundary parameter
2020 Mathematics Subject Classification:
35P15, 58J50, 52A55
The authors were supported by NSFC grant No. 11831005 and NSFC-FWO 11961131001.
1. Introduction
Let $\Omega$ be a bounded domain with smooth boundary in an $n$-dimensional complete Riemannian manifold $\left(M,g\right)$. The following is called the eigenvalue problem of the Robin Laplacian
(1.1) $\begin{cases}-\Delta u=\lambda\left(\Omega,\beta\right)~u,&\text{in}~\Omega,\\ \dfrac{\partial u}{\partial N}+\beta u=0,&\text{on }\partial\Omega,\end{cases}$
where $N$ is the unit outward normal to $\partial\Omega$ and $\beta\in\mathbb{R}$. For the Robin eigenvalue problem (1.1), the associated Rayleigh quotient is
$\mathcal{R}(u)=\dfrac{\displaystyle\int_{\Omega}|\nabla u|^{2}\,d\mu+\beta\int_{\partial\Omega}u^{2}\,d\sigma}{\displaystyle\int_{\Omega}u^{2}\,d\mu},\quad u\in H^{1}\left(\Omega\right)\setminus\{0\}.$
The first eigenvalue of the Robin Laplacian is characterized by the variational principle
(1.2) $\lambda_{1}\left(\Omega,\beta\right)=\min_{\begin{subarray}{c}u\in H^{1}(\Omega)\\ u\neq 0\end{subarray}}\mathcal{R}(u).$
When $\beta<0$, Bareket [Bar77] conjectured in 1977 that the ball maximizes $\lambda_{1}\left(\Omega,\beta\right)$ among all Lipschitz sets of a given area in the Euclidean plane. In 2015, Ferone, Nitsch and Trombetti [FNT15] established that the ball is a local maximizer for the first eigenvalue of the Robin Laplacian within the class of bounded Lipschitz domains in $\mathbb{R}^{n}$. Freitas and Krejčiřík [FK15] showed that the Bareket conjecture holds for $\beta$ sufficiently close to 0 in the Euclidean plane. For a compact Riemannian manifold, Savo [Sav20] established an upper bound for the first eigenvalue of the Robin Laplacian in terms of the dimension, inradius, and lower bounds on Ricci and boundary mean curvature. Later, Li and Wang [LW21] proved an upper bound for the first eigenvalue of the $p$-Laplacian with Robin boundary conditions in terms of these geometric parameters, via comparison with an associated one-dimensional eigenvalue problem. Li, Wang and Wu [LWW23] established that geodesic balls in nonpositively curved space forms maximize the second eigenvalue of the Robin Laplacian among bounded domains with the same volume.
For bounded planar domains $\Omega$ of class $C^{2}$ in $\mathbb{R}^{2}$, Antunes, Freitas and Krejčiřík [AFK17] proved that
(1.3) $\lambda_{1}\left(\Omega,\beta\right)\leq\lambda_{1}\left(\Omega^{\star},\beta\right),$
where $\Omega^{\star}$ is a ball with the same perimeter as $\Omega$. Bucur et al. [BFN+19] showed that (1.3) holds for convex domains in $\mathbb{R}^{n}$. Amato, Gentile and Masiello [AGM24] established a lower bound, in terms of the isoperimetric deficit, for the first eigenvalue of the $p$-Laplacian with Robin boundary conditions on convex sets with prescribed perimeter in $\mathbb{R}^{n}$. For more on eigenvalue estimates, see [FL21, KL22, Vik22], among others.
In [ACC+25], Acampora et al. established a lower bound for the first eigenvalue of the Robin Laplacian with negative parameter on strongly convex sets in the sphere $\mathbb{S}^{n}$, under a perimeter constraint. The authors [ACC+25] also constructed convex sets for which the inner parallel sets are not convex in hyperbolic space $\mathbb{H}^{n}$. Motivated by [ACC+25], we observe that the inner parallel sets of horospherically convex bounded domains are also horospherically convex in $\mathbb{H}^{n}$. Based on this observation, in the first part, for $\beta<0$, we establish a lower bound for $\lambda_{1}\left(\Omega,\beta\right)$ on horospherically convex bounded domains in the hyperbolic space $\mathbb{H}^{n}$ by means of the isoperimetric deficit.
Theorem 1.1.
Let $\Omega$ be a horospherically convex bounded domain with smooth boundary in hyperbolic space $\mathbb{H}^{n}$ and $P(\Omega)$ be the perimeter of $\Omega$. Let $\Omega^{\star}$ be the geodesic ball with the same perimeter as $\Omega$, that is, $P\left(\Omega\right)=P\left(\Omega^{\star}\right)$. For $\beta<0$, denote by $v$ the eigenfunction associated with $\lambda_{1}(\Omega^{\star},\beta)$ for (1.1). Then
(1.4) $\frac{\lambda_{1}(\Omega^{\star},\beta)-\lambda_{1}(\Omega,\beta)}{|\lambda_{1}(\Omega,\beta)|}\geq\frac{v_{m}^{2}}{\|v\|_{L^{2}(\Omega^{\star})}^{2}}\bigl(|\Omega^{\star}|-|\Omega|\bigr),$
where $v_{m}=\min\limits_{x\in\Omega^{\star}}v(x)$ and $\left|\Omega\right|$ is the volume of $\Omega$.
According to the isoperimetric inequality in hyperbolic space [Cha84], the following result holds.
Corollary 1.1.
Under the same assumptions as in Theorem 1.1, we have
(1.5) $\lambda_{1}\left(\Omega,\beta\right)\leq\lambda_{1}\left(\Omega^{\star},\beta\right).$
Moreover, equality holds if and only if $\Omega$ is isometric to a geodesic ball in $\mathbb{H}^{n}$.
Remark 1.1.
Bucur et al. [BFN+19] established (1.5) for convex domains in $\mathbb{R}^{n}$. Acampora et al. [ACC+25] proved that both (1.4) and (1.5) hold on strongly convex domains in $\mathbb{S}^{n}$.
For the Robin eigenvalue problem (1.1) with $\beta>0$, Bossel [Bos86] and Daners [Dan06] proved that, in Euclidean space,
(1.6) $\lambda_{1}(\Omega,\beta)\geq\lambda_{1}\left(\Omega^{\sharp},\beta\right),$
where $\Omega^{\sharp}$ denotes an Euclidean ball with the same volume as $\Omega$. Bucur and Daners [BD10], Dai and Fu [DF11] independently proved the inequality (1.6) for the $p$-Laplacian in Euclidean space. The Bossel-Daners inequality was established for complete noncompact Riemannian manifolds with $\mathrm{Ric}\geq 0$ in [CLW23], $n$-dimensional compact Riemannian manifolds with $\mathrm{Ric}\geq n-1$ and $n$-dimensional hyperbolic space in [CCL22]. Furthermore, for convex sets with prescribed perimeter in $\mathbb{R}^{n}$, Amato, Gentile and Masiello [AGM24] obtained an upper bound for the first eigenvalue of the $p$-Laplacian with Robin boundary conditions in terms of the isoperimetric deficit. Analogous results can be found in [DK08, CG22, PP24] and references therein.
In the second part, we focus on the estimates for the first eigenvalue of the Robin Laplacian with positive boundary parameter in hyperbolic space $\mathbb{H}^{n}$. For horospherically convex bounded domains with smooth boundary in hyperbolic space $\mathbb{H}^{n}$, we establish an upper bound for the first eigenvalue of the Robin Laplacian with positive boundary parameter.
Theorem 1.2.
Let $\Omega$ be a horospherically convex bounded domain with smooth boundary in hyperbolic space $\mathbb{H}^{n}$ and $\Omega^{\star}$ be the geodesic ball with the same perimeter as $\Omega$, that is, $P\left(\Omega\right)=P\left(\Omega^{\star}\right)$. For $\beta>0$, denote by $v$ the eigenfunction associated with $\lambda_{1}(\Omega^{\star},\beta)$ for (1.1). Then
(1.7) $\frac{\lambda_{1}(\Omega,\beta)-\lambda_{1}(\Omega^{\star},\beta)}{\lambda_{1}(\Omega,\beta)}\leq\frac{v_{M}^{2}}{\|v\|_{L^{2}(\Omega^{\star})}^{2}}\left(|\Omega^{\star}|-|\Omega|\right),$
where $v_{M}=\max\limits_{x\in\Omega^{\star}}v(x)$.
Remark 1.2.
For $\beta=\infty$, equation (1.1) can be seen as the problem equipped with the Dirichlet boundary condition; the inequality in Theorem 1.2 provides an estimate for the Faber-Krahn deficit. In this case, Brandolini, Nitsch and Trombetti [BNT10] derived a Faber-Krahn deficit in terms of the isoperimetric deficit in $\mathbb{R}^{n}$.
Remark 1.3.
Let $\Omega\subset\mathbb{S}^{n}$ be an open set such that $\overline{\Omega}$ is strongly convex and $\Omega^{\star}$ be the strongly convex geodesic ball with the same perimeter as $\Omega$. For $\beta>0$, it holds that
(1.8) $\frac{\lambda_{1}(\Omega,\beta)-\lambda_{1}(\Omega^{\star},\beta)}{\lambda_{1}(\Omega,\beta)}\leq\frac{v_{M}^{2}}{\|v\|_{L^{2}(\Omega^{\star})}^{2}}\left(|\Omega^{\star}|-|\Omega|\right).$
Combining the isoperimetric inequality [Cha84], the Bossel-Daners inequality [CCL22, CLW23], and the monotonicity of the first eigenvalue of the Robin Laplacian on concentric balls, we conclude that the left-hand side of each of (1.7) and (1.8) is nonnegative.
The paper is organized as follows. In Section 2, we recall some notions, the Steiner formula and the Alexandrov-Fenchel inequality in hyperbolic space. This section also analyzes the convexity of the parallel domains and provides an estimate for the perimeter of the inner parallel domains. In the final section, we present complete proofs of Theorems 1.1 and 1.2.
2. Preliminaries
This section recalls some basic notions and properties in hyperbolic space. For details, we refer to [Koh91, WX14]. Furthermore, we analyze the convexity of parallel domains and derive an estimate for the perimeter of the inner parallel domains.
Let $\mathbb{R}_{1}^{n+1}$ be the real vector space $\mathbb{R}^{n+1}$ equipped with the Lorentzian metric
$\langle{x},{y}\rangle=\langle{\left(x_{0},x_{1},\cdots,x_{n}\right)},{\left(y_{0},y_{1},\cdots,y_{n}\right)}\rangle=-x_{0}y_{0}+x_{1}y_{1}+\dots+x_{n}y_{n}.$
The real hyperbolic space with constant sectional curvature 1-1 is given by
$\mathbb{H}^{n}=\left\{x\in\mathbb{R}_{1}^{n+1}\big|~\langle{x},{x}\rangle=-1,x_{0}\geq 1\right\}.$
2.A. Quermassintegrals
Let $K\subset\mathbb{H}^{n}$ be a convex domain with smooth boundary. For every $t\geq 0$, define the inner parallel set $K_{t}$ and the outer parallel set $K^{t}$ of $K$ as follows
$K_{t}=\left\{x\in K:d\left(x,\,\partial K\right)\geq t\right\},\qquad K^{t}=\left\{x\in\mathbb{H}^{n}:d\left(x,K\right)\leq t\right\},$
where $d\left(\cdot,\cdot\right)$ is the Riemannian distance, $d\left(x,\,\partial K\right)$ is the distance from $x\in K$ to the boundary $\partial K$ and $d(x,K)=\inf\limits_{q\in K}d(x,q)$. For every $p\in\partial K$, we denote by $H_{k}$ the $k$-th normalized symmetric function of the principal curvatures $\kappa_{1},\ldots,\kappa_{n-1}$ of $\partial K$,
$H_{k}=\frac{\sum\limits_{1\leq i_{1}<\ldots<i_{k}\leq n-1}\kappa_{i_{1}}\cdots\kappa_{i_{k}}}{\binom{n-1}{k}},\quad k=1,\ldots,n-1,$
and $H_{0}=1$. In particular, $H_{1}$ is defined as the mean curvature of $\partial K$ at $p$. For some positive number $s>0$, the Steiner formula for the perimeter reads as
$P\left(K^{s}\right)=\sum_{i=0}^{n-1}\binom{n-1}{i}\cosh^{i}s\,\sinh^{n-1-i}s\,V_{i}\left(K\right),$
where $V_{i}(K)$ are curvature integrals on $K$ defined by
$V_{i}(K)=\int_{\partial K}H_{n-1-i}\,d\mu,\quad i=0,\ldots,n-1.$
Then
(2.1) $\frac{d}{ds}P\left(K^{s}\right)\biggl|_{s=0}=(n-1)V_{n-2}\left(K\right).$
The quermassintegrals of $K$ are defined as follows
$W_{k}(K):=\frac{(n-k)\omega_{k-1}\cdots\omega_{0}}{n\omega_{n-2}\cdots\omega_{n-k-1}}\int_{\mathcal{L}_{k}}\chi(L_{k}\cap K)\,dL_{k},\quad k=1,\ldots,n-1,$
where $\omega_{k}$ denotes the Hausdorff measure of the $k$-dimensional unit sphere, $\mathcal{L}_{k}$ is the space of $k$-dimensional totally geodesic subspaces $L_{k}$ in $\mathbb{H}^{n}$, $dL_{k}$ is the natural (invariant) measure on $\mathcal{L}_{k}$ and $\chi$ is the characteristic function of $K$. In particular,
$W_{0}(K)=\left|K\right|,\quad W_{1}(K)=\frac{1}{n}P\left(K\right),\quad W_{n}(K)=\frac{\omega_{n-1}}{n}.$
From Proposition 7 in [Sol05], the quermassintegrals and the curvature integrals are related by
$V_{n-1-k}(K)=n\left(W_{k+1}(K)+\frac{k}{n-k+1}W_{k-1}(K)\right),\quad k=1,\ldots,n-1,$
and
$V_{n-1}(K)=nW_{1}(K)=P\left(K\right).$
2.B. Parallel hypersurfaces and Alexandrov-Fenchel inequality
For a domain $\Omega\subset\mathbb{H}^{n}$, denote by $R_{\Omega}$ the inradius of $\Omega$, that is,
$R_{\Omega}=\max_{x\in\Omega}d\left(x,\,\partial\Omega\right).$
Recall the following proposition, which may be known to experts.
Proposition 2.1.
Let $\Omega$ be a horospherically convex bounded domain with smooth boundary $\partial\Omega$ in $\mathbb{H}^{n}$. Then the set $\Omega_{t}$, $0\leq t<R_{\Omega}$, is a horospherically convex domain, and $\Omega^{s}$ is also a horospherically convex domain for $s>0$.
Proof.
For readers’ convenience, we provide a proof of the proposition by adapting the argument of [MR91]. Since $\mathbb{H}^{n}\subset\mathbb{R}_{1}^{n+1}$, let $x:\partial\Omega\rightarrow\mathbb{R}_{1}^{n+1}$ be a map with $\langle{x},{x}\rangle=-1$ and $x_{0}\geq 1$. Denote by $\exp$ the exponential map of $\mathbb{H}^{n}$ and by $N$ the unit outward normal vector field to $\partial\Omega$. For the set $\Omega_{t}$, define
$x_{t}(p)=\exp_{x(p)}\left(-tN(p)\right)=\cosh t\,x(p)-\sinh t~N(p),\qquad p\in\partial\Omega.$
Let $\{\kappa_{i}\}_{i=1}^{n-1}$ be the principal curvatures of $\partial\Omega$ at $p$, with corresponding orthonormal principal directions $\{e_{i}\}_{i=1}^{n-1}$. Then,
$(x_{t})_{*}(e_{i})=\left(\cosh t-\kappa_{i}\sinh t\right)e_{i},$
and hence, $N_{t}=-\sinh t\,x+\cosh t\,N$ is a unit normal field for $x_{t}$. Moreover, the second fundamental form $\sigma_{t}$ of $x_{t}$ is given by
$\sigma_{t}\left((x_{t})_{*}(e_{i}),(x_{t})_{*}(e_{j})\right)=\langle{(N_{t})_{*}(e_{i})},{(x_{t})_{*}(e_{j})}\rangle.$
By direct computation,
$\kappa_{i}(t)=\frac{-\sinh t+\kappa_{i}\cosh t}{\cosh t-\kappa_{i}\sinh t}=\frac{\kappa_{i}-\tanh t}{1-\kappa_{i}\tanh t}.$
Then
$\kappa_{i}(t)-1=\frac{\left(1+\tanh t\right)\left(\kappa_{i}-1\right)}{1-\kappa_{i}\tanh t}\geq 0.$
Hence, $\Omega_{t}$ is a horospherically convex domain.
For the set $\Omega^{s}$, $s>0$, we can adopt a similar argument to that for $\Omega_{t}$ and conclude that $\Omega^{s}$ is a horospherically convex domain. ∎
For any horospherically convex bounded domain with smooth boundary in $\mathbb{H}^{n}$, the following hyperbolic Alexandrov-Fenchel inequality holds.
Theorem 2.1.
[WX14, Theorem 1.3] Let $1\leq k\leq n-1$. Any horospherically convex bounded domain $\Omega$ with smooth boundary in $\mathbb{H}^{n}$ satisfies
$\int_{\partial\Omega}H_{k}\,d\sigma\geq\omega_{n-1}\left\{\left(\frac{P\left(\Omega\right)}{\omega_{n-1}}\right)^{\frac{2}{k}}+\left(\frac{P\left(\Omega\right)}{\omega_{n-1}}\right)^{\frac{2}{k}\frac{(n-k-1)}{n-1}}\right\}^{\frac{k}{2}}.$
Equality holds if and only if $\Omega$ is a geodesic ball.
By combining Proposition 2.1 with Theorem 2.1, we obtain an estimate for the derivative of the perimeter of $\Omega_{t}$.
Lemma 2.1.
Let $\Omega$ be a horospherically convex bounded domain with smooth boundary in $\mathbb{H}^{n}$. Then for almost every $t\in\left(0,\,R_{\Omega}\right)$,
$\frac{d}{dt}P\left(\Omega_{t}\right)\geq(n-1)\omega_{n-1}\left\{\left(\frac{P\left(\Omega_{t}\right)}{\omega_{n-1}}\right)^{2}+\left(\frac{P\left(\Omega_{t}\right)}{\omega_{n-1}}\right)^{\frac{2(n-2)}{n-1}}\right\}^{\frac{1}{2}}.$
Equality holds if $\Omega$ is a geodesic ball.
Proof.
By Proposition 2.1, the set $\Omega_{t}$ is a horospherically convex domain, which implies that $\left(\Omega_{t}\right)^{s}$ is also a horospherically convex domain for $s>0$. Notice that
$\left(\Omega_{t}\right)^{s}\subset\Omega_{t-s}.$
The function $P\left(\Omega_{t}\right)$ is decreasing with respect to $t$, hence it is differentiable almost everywhere. For almost every $t\in\left(0,R_{\Omega}\right)$, due to (2.1), one can infer that
$-\frac{d}{dt}P\left(\Omega_{t}\right)$ $=\lim_{s\to 0^{+}}\frac{P(\Omega_{t-s})-P(\Omega_{t})}{s}$
$\geq\lim_{s\to 0^{+}}\frac{P((\Omega_{t})^{s})-P(\Omega_{t})}{s}$
$=(n-1)V_{n-2}\left(\Omega_{t}\right).$
By the definition of $V_{n-2}\left(\Omega_{t}\right)$ and the hyperbolic Alexandrov-Fenchel inequality in Theorem 2.1 with $k=1$, the proof of Lemma 2.1 is complete. ∎
3. Proofs of Theorem 1.1 and 1.2
In this section, inspired by [Pol60, PW61, ACC+25], we prove Theorems 1.1 and 1.2 using the method of parallel coordinates to construct a suitable test function.
Without loss of generality, we assume that $\Omega^{\star}$ is the geodesic ball centered at $q\in\mathbb{H}^{n}$ with radius $R>0$ satisfying $P\left(\Omega^{\star}\right)=P\left(\Omega\right)$. The eigenfunction $v(x)=\psi\left(d\left(x,q\right)\right)>0$ associated with $\lambda_{1}\left(\Omega^{\star},\beta\right)$ is the radial solution to the following equation
(3.1) $\begin{cases}\psi^{\prime\prime}+(n-1)\dfrac{\cosh r}{\sinh r}\psi^{\prime}+\lambda_{1}\left(\Omega^{\star},\beta\right)\psi=0&r\in(0,R),\\ \psi^{\prime}(0)=0,\\ \psi^{\prime}(R)+\beta\psi(R)=0.\end{cases}$
Letting $\phi\left(\rho\right)=\psi\left(R-\rho\right)$, we can regard $v$ as a function of the distance to the boundary of the ball $\Omega^{\star}$, that is,
$v(x)=\phi\left(d\left(x,\partial\Omega^{\star}\right)\right).$
Choosing $u=1$ in (1.2), it can be concluded that $\lambda_{1}\left(\Omega^{\star},\beta\right)<0$ for $\beta<0$. Moreover, by (3.1),
$\dfrac{\partial}{\partial r}\left(\sinh^{n-1}r\,v^{\prime}(r)\right)=-\lambda_{1}\left(\Omega^{\star},\beta\right)v\left(r\right)\sinh^{n-1}r>0,$
which implies that $v^{\prime}(r)>0$ for $r\in(0,R]$, that is, $v\left(r\right)$ is strictly increasing with respect to $r$. Similarly, if $\beta>0$, then $v^{\prime}(r)<0$ for $r\in(0,R]$, that is, $v\left(r\right)$ is strictly decreasing with respect to $r$.
Proof of Theorem 1.1.
Let $v>0$ be the eigenfunction associated with the eigenvalue $\lambda_{1}\left(\Omega^{\star},\beta\right)$ for (3.1). Since $v^{\prime}(r)>0$ for $r\in(0,R]$, where $R$ is the radius of $\Omega^{\star}$, one has $\phi^{\prime}(t)<0$ for $t\in[0,R)$. Define
$u(x)=\phi\left(d\left(x,\,\partial\Omega\right)\right),\qquad x\in\Omega.$
Then
(3.2) $v_{m}=\min\limits_{x\in\Omega^{\star}}v(x)=\phi\left(R\right)\leq\phi\left(R_{\Omega}\right)=\min\limits_{x\in\Omega}u(x)=u_{m}.$
By Lemma 2.1, we have
$\frac{d}{dt}P\left(\Omega_{t}\right)\leq-\left(n-1\right)\omega_{n-1}\left\{\left(\frac{P\left(\Omega_{t}\right)}{\omega_{n-1}}\right)^{2}+\left(\frac{P\left(\Omega_{t}\right)}{\omega_{n-1}}\right)^{\frac{2(n-2)}{n-1}}\right\}^{\frac{1}{2}}$
and
$\frac{d}{dt}P\left(\Omega^{\star}_{t}\right)=-\left(n-1\right)\omega_{n-1}\left\{\left(\frac{P\left(\Omega^{\star}_{t}\right)}{\omega_{n-1}}\right)^{2}+\left(\frac{P\left(\Omega^{\star}_{t}\right)}{\omega_{n-1}}\right)^{\frac{2(n-2)}{n-1}}\right\}^{\frac{1}{2}}.$
From Lemma 3.3 in [ACC+25], it follows that
(3.3) $P\left(\Omega_{t}\right)\leq P\left(\Omega^{\star}_{t}\right),\qquad t\in\left(0,R_{\Omega}\right).$
Combining (3.2), (3.3) with the co-area formula yields
$\begin{split}\int_{\Omega}u^{2}\,d\mu&=\int_{0}^{R_{\Omega}}\phi^{2}(t)P\left(\Omega_{t}\right)\,dt\\ &=\int_{0}^{R}\phi^{2}(t)P\left(\Omega^{\star}_{t}\right)\,dt-\int_{0}^{R}\phi^{2}(t)\left(P\left(\Omega^{\star}_{t}\right)-P\left(\Omega_{t}\right)\right)\,dt\\ &\leq\int_{\Omega^{\star}}v^{2}\,d\mu-v_{m}^{2}\left(\left|\Omega^{\star}\right|-\left|\Omega\right|\right),\end{split}$
and
$\int_{\Omega}|\nabla u|^{2}\,d\mu=\int_{0}^{R_{\Omega}}(\phi^{\prime}(t))^{2}P(\Omega_{t})\,dt\leq\int_{0}^{R}(\phi^{\prime}(t))^{2}P(\Omega^{\star}_{t})\,dt=\int_{\Omega^{\star}}|\nabla v|^{2}\,d\mu.$
A direct computation shows that
$\int_{\partial\Omega}u^{2}\,d\sigma=\phi^{2}(0)P\left(\Omega\right)=\int_{\partial\Omega^{\star}}v^{2}\,d\sigma.$
Hence, we have
$\lambda_{1}\left(\Omega,\beta\right)$ $\leq\dfrac{\displaystyle\int_{\Omega}|\nabla u|^{2}\,d\mu+\beta\int_{\partial\Omega}u^{2}\,d\sigma}{\displaystyle\int_{\Omega}u^{2}\,d\mu}$
$\leq\dfrac{\displaystyle\int_{\Omega^{\star}}|\nabla v|^{2}\,d\mu+\beta\int_{\partial\Omega^{\star}}v^{2}\,d\sigma}{\displaystyle\int_{\Omega^{\star}}v^{2}\,d\mu-v_{m}^{2}\left(\left|\Omega^{\star}\right|-\left|\Omega\right|\right)}$
$=\dfrac{\displaystyle\lambda_{1}\left(\Omega^{\star},\beta\right)}{\displaystyle 1-\frac{\displaystyle v^{2}_{m}\left(\left|\Omega^{\star}\right|-\left|\Omega\right|\right)}{\displaystyle\|v\|_{L^{2}\left(\Omega^{\star}\right)}^{2}}}.$
This completes the proof of Theorem 1.1. ∎
Proof of Theorem 1.2.
For $\beta>0$, let $v>0$ be the eigenfunction associated with $\lambda_{1}\left(\Omega^{\star},\beta\right)$ for the problem (3.1). Then $\phi^{\prime}(t)>0$ for $t\in[0,R)$. Define
$u(x)=\phi\left(d\left(x,\,\partial\Omega\right)\right).$
Then
(3.4) $u_{M}=\max\limits_{x\in\Omega}u(x)=\phi\left(R_{\Omega}\right)\leq\phi\left(R\right)=\max\limits_{x\in\Omega^{\star}}v(x)=v_{M}.$
The definitions of $u$ and $v$ yield
(3.5) $\int_{\partial\Omega}u^{2}\,d\sigma=\phi^{2}(0)P\left(\Omega\right)=\phi^{2}(0)P\left(\Omega^{\star}\right)=\int_{{\partial\Omega}^{\star}}v^{2}\,d\sigma.$
From (3.3), (3.4) and the co-area formula, it follows that
(3.6) $\begin{split}\int_{\Omega}\left|\nabla u\right|^{2}\,d\mu&=\int_{0}^{R_{\Omega}}\left(\phi^{\prime}(t)\right)^{2}P\left(\Omega_{t}\right)\,dt\\ &\leq\int_{0}^{R}\left(\phi^{\prime}(t)\right)^{2}P\left(\Omega^{\star}_{t}\right)\,dt\\ &=\int_{\Omega^{\star}}\left|\nabla v\right|^{2}\,d\mu,\end{split}$
and
(3.7) $\begin{split}\int_{\Omega}u^{2}\,d\mu=&\int_{0}^{R_{\Omega}}\phi^{2}(t)P\left(\Omega_{t}\right)\,dt\\ =&\int_{0}^{R}\phi^{2}(t)P\left(\Omega^{\star}_{t}\right)\,dt-\int_{0}^{R}\phi^{2}(t)\left(P\left(\Omega^{\star}_{t}\right)-P\left(\Omega_{t}\right)\right)\,dt\\ \geq&\int_{\Omega^{\star}}v^{2}\,d\mu-v_{M}^{2}\left(\left|\Omega^{\star}\right|-\left|\Omega\right|\right).\end{split}$
If the inequality
$\int_{\Omega^{\star}}v^{2}\,d\mu-v_{M}^{2}\left(|\Omega^{\star}|-|\Omega|\right)\leq 0$
holds, it follows that
$\frac{\displaystyle\lambda_{1}(\Omega,\beta)-\lambda_{1}(\Omega^{\star},\beta)}{\displaystyle\lambda_{1}(\Omega,\beta)}\leq 1\leq\frac{\displaystyle v_{M}^{2}\left(|\Omega^{\star}|-|\Omega|\right)}{\displaystyle\|v\|_{L^{2}(\Omega^{\star})}^{2}}.$
Otherwise, if
$\int_{\Omega^{\star}}v^{2}\,d\mu-v_{M}^{2}\left(|\Omega^{\star}|-|\Omega|\right)>0,$
then combining (3.5), (3.6) and (3.7) yields
$\lambda_{1}\left(\Omega,\beta\right)$ $\leq\frac{\displaystyle\int_{\Omega}|\nabla u|^{2}\,d\mu+\beta\int_{\partial\Omega}u^{2}\,d\sigma}{\displaystyle\int_{\Omega}u^{2}\,d\mu}$
$\leq\frac{\displaystyle\int_{\Omega^{\star}}|\nabla v|^{2}\,d\mu+\beta\int_{\partial\Omega^{\star}}v^{2}\,d\sigma}{\displaystyle\int_{\Omega^{\star}}v^{2}\,d\mu-v_{M}^{2}\left(|\Omega^{\star}|-|\Omega|\right)}$
$=\frac{\lambda_{1}(\Omega^{\star},\beta)}{\displaystyle 1-\frac{v_{M}^{2}\left(|\Omega^{\star}|-|\Omega|\right)}{\|v\|_{L^{2}(\Omega^{\star})}^{2}}}.$
This completes the proof of Theorem  1.2. ∎